OpenStudy (adunb8):

differential equation & linear algebra equation. I need to know how to use Bernoulli Equation to solve for xy' + y + x^2y^2e^x

OpenStudy (adunb8):

i want to know the step by step.

OpenStudy (anonymous):

\[xy'+y=x^2e^xy^2\]

OpenStudy (adunb8):

uh... step by steps?

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Bernoulli_differential_equation \[z=y^{1-2}=y^{-1}\]

OpenStudy (adunb8):

.... okay thanks

OpenStudy (lgbasallote):

lol..wikipedia

OpenStudy (adunb8):

oh! can you help?!

OpenStudy (lgbasallote):

i'll let @mukushla do his thing

OpenStudy (adunb8):

....

OpenStudy (adunb8):

his using wikipedia...

OpenStudy (anonymous):

\[z'=-\frac{y'}{y^2}\]\[xy'+y=x^2e^xy^2\]divide by y^2\[x\frac{y'}{y^2}+\frac{1}{y}=x^2e^x\]now u have\[-xz'+z=x^2e^x\]

OpenStudy (anonymous):

this is a first order linear diff equation solve it for z

OpenStudy (adunb8):

okay thank you for the help ~.~

OpenStudy (anonymous):

welcome

OpenStudy (lgbasallote):

do you know the form of Bernoulli's equation @adunb8 ?

OpenStudy (adunb8):

well know i looked at it it looks like its \[y' + p(x)y = q(x)y^n\]

OpenStudy (lgbasallote):

right

OpenStudy (lgbasallote):

so your first step should be to transform your equation into that form

OpenStudy (adunb8):

yes i made it into \[y' + x^{-1} y = xy^2e^x\]

OpenStudy (lgbasallote):

wait....what's the original equation?

OpenStudy (adunb8):

\[xy' + y + x ^{2}y ^{2}e ^{x}\]

OpenStudy (adunb8):

= 0